OpenStudy (strawberry17):

How can (2√13)/13 be simplified to 2/√13 ?

OpenStudy (compassionate):

Can you write it using the equation editor.

OpenStudy (strawberry17):

\[ \frac{2\sqrt{13} }{ 13 } \to \frac{ 2 }{\sqrt{13} ? }\]

OpenStudy (strawberry17):

I know they both equal the same thing when calculated..

OpenStudy (compassionate):

Gee. I'm actually stumped on this one

OpenStudy (strawberry17):

Same here, at first I thought the answer manual made a mistake, until I checked to see that they actually equal the same thing

OpenStudy (compassionate):

You broke math.

OpenStudy (strawberry17):

lol

OpenStudy (jdoe0001):

\(\bf \cfrac{2\sqrt{13} }{ 13 } \times \cfrac{\sqrt{13}}{\sqrt{13}} \implies \cfrac{2\sqrt{13}\sqrt{13}}{13\sqrt{13}} \implies \cfrac{2\sqrt{13^2}}{13\sqrt{13}} \implies \cfrac{2\times \cancel{13}}{\cancel{13}\sqrt{13}}\)

OpenStudy (strawberry17):

Thanks! That really explains it well :)

OpenStudy (compassionate):

It works swell when you do it that way. But you're changing the value by adding this things.

OpenStudy (compassionate):

Alle von dem

OpenStudy (jdoe0001):

well, keep in mind that \(\bf \cfrac{\sqrt{13}}{\sqrt{13}} = 1\) so all I really did was multiply by 1, which isn't chaning the value

OpenStudy (jdoe0001):

1,000,000 times 1 = 1,000,000

OpenStudy (compassionate):

But you're putting those things there to cancel them, If it equals one, then do the same thing with just 1, and not use the 13/13/. You'll find that it doesn't work. I'm saying that this is a fault in mathematics, things like this, because they're not consistent with their exact equals. So it doesn't work. We just accept these things as true and move on.

OpenStudy (jdoe0001):

heheh, well, it doesn't change the value, one can say that \(\bf \cfrac{2\sqrt{13} }{ 13 }\) was really obtained the same way from \(\bf \cfrac{2}{\sqrt{13}}\) btw \(\bf \cfrac{2\sqrt{13} }{ 13 } = 0.554700196\\ \bf \cfrac{2}{\sqrt{13}} = 0.554700196\)

OpenStudy (jdoe0001):

there's only 1 discrepancy in the whole thing, which is fairly understood when using it, and is when the domain is relevant, but in this case it isn't